All answers must be entered as formulas (using cell names)Y3K, Inc., has sales of $5,783, total assets of $2,604, and a debt-equity ratio of .75. If its return on equity is 11 percent, what is its net income? 4 Total assets Sales Debt-equity ratio Return on equity 2,604 5,783 0.75 11% 10 Complete the following analysis. Do not hard code values in your calculations Profit margin Net income 12 14 15 16 19 20
Answer #1
Calculation of net income:
Let equity be "x"
then debt= 0.75*equity= 0.75x
Total assets= Debt+equity
2604= x+0.75x
2604= 1.75x
x= 2604/1.75= 1488
Equity= $1488
Net income= Return on equity*equity
Net income= 0.11*1488= $163.68
Calculation of profit margin:
Profit margin= net income/ sales*100
Profit margin= 163.68/5783*100= 2.83%
Profit margin= 2.83%